# Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance

### v2

@@ Abstract
While large language models (LLMs) have shown strong general reasoning capabilities, their effectiveness in financial reasoning—which is crucial for real-world financial applications—remains underexplored. In this study, we conduct a comprehensive evaluation of 24 state-of-the-art general and reasoning-focused LLMs across four complex financial reasoning tasks involving financial text, tabular data, and equations. We assess key capabilities such as numerical reasoning, tabular interpretation, financial terminology comprehension, long-context understanding, and equation-based problem solving. Our analysis reveals that while data quality and pretraining contribute to performance, general techniques like chain-of-thought (CoT) fine-tuning offer limited gains in financial tasks. To address this, we propose two domain-adapted models, Fino1-8B and Fino1-14B, trained with CoT fine-tuning and reinforcement learning using domain-specific reasoning paths. Our models are trained on a carefully curated dataset integrating high-quality examples from diverse sources, covering financial reports, tables, equations, and structured XBRL texts. Despite limited training data, they achieve an 7–9% performance improvement, outperforming several advanced LLMs—including GPT-o1, GPT-o3-mini, GPT-4.5, and comparable with DeepSeek models (V3 and R1)—demonstrating strong practical value in resource-constrained scenarios. Our findings highlight the need for domain-specific adaptations in financial reasoning, and we release all datasets, models, and code for future research [^1] https://github.com/The-FinAI/Fino1.
@@ 1 Introduction
Advancements in large language models (LLMs) have demonstrated remarkable performance across various natural language pocessing tasks, including content generationVan Veen et al. (2024); Zhang et al. (2024), language translation Xu et al. (2024), and sentiment analysis Xing (2024); Miah et al. (2024). More recently, reasoning-enhanced models such as OpenAI’s o1 and DeepSeek’s R1 have been developed to extend LLM capabilities in complex reasoning Guo et al. (2025); Jaech et al. (2024). These models exhibit significant improvements, particularly in mathematical and logical tasks that require complex reasoning Temsah et al. (2024); Zhong et al. (2024). However, despite their success in general reasoning, their performance in the financial domain remains largely unexplored.
Financial reasoning is essential in the financial domain, where professionals must extract key insights from lengthy reports and multi-table documents. It requires not only numerical computation but also a deep understanding of financial terminology, regulations, and economic concepts Xie et al. (2023, 2024a). Unlike general reasoning tasks, financial reasoning involves (1) domain-specific language comprehension, (2) reasoning over financial concepts and numbers, and (3) interpreting structured data such as financial tables Xie et al. (2024b). Evaluating LLMs in this context is critical for closing the gap between general AI reasoning and real-world financial applications.
To address this gap, we firstly comprehensively evaluate the performance of existing powerful reasoning models and conduct an in-depth analysis of their capabilities in financial tasks, to provide insights into the strengths and limitations of existing reasoning models in financial applications. We focus on the following key aspects: (1) Transferability of general-domain reasoning enhancements – Assessing whether reasoning-enhanced LLMs, originally optimized for general-domain reasoning, also demonstrate improved performance in financial reasoning tasks. (2) Impact across financial tasks – Analyzing how the reasoning enhancement process influences LLM performance across different types of financial tasks. (3) Scaling and performance gap – Investigating the differences between smaller reasoning models and large-scale LLMs in financial reasoning. (4) Limits of LLMs in complex financial tasks – Evaluating the extent to which LLMs can handle extremely complex financial reasoning challenges.
We evaluated 24 LLMs from the GPT Achiam et al. (2023), LLaMA Dubey et al. (2024), DeepSeek Liu et al. (2024), and Qwen Yang et al. (2024) families on four financial reasoning benchmarks: FinQA Chen et al. (2021), DM-Simplong Zhao et al. (2024a), XBRL-Math Wang et al. (2025) and DM-Complong. These datasets span financial texts, tables, and equations, enabling evaluation of models in numerical reasoning, domain-specific terminology comprehension, long-context processing, and equation-based reasoning tasks. Our evaluation encompasses both general-purpose and reasoning-enhanced models, ranging from smaller-scale models (e.g., LLaMA 3-8B-InstructDubey et al. (2024)) to high-capacity models (e.g., GPT-o1Jaech et al. (2024)), as well as reasoning-augmented models incorporating strategies such as data sampling and financial Chain-of-Thought (CoT) construction. Experimental results reveal three key insights: (1) Reasoning-enhanced models like DeepSeek-R1 and GPT-o1 do not consistently outperform general-purpose models on financial tasks, likely due to challenges related to domain-specific terminology and multi-table reasoning; (2) Performance gains from larger model sizes and improved pretraining tend to plateau beyond 32B or 70B parameters;
To develop reasoning models tailored for financial domain, we build novel financial reasoning-enhanced large language models Fino1-8B and Fino1-14B based on LLaMA-3.1-8B-InstructDubey et al. (2024) and Qwen2.5-14B-InstructYang et al. (2024), chosen for their strong performance among smaller-scale LLMs. To enhance reasoning capabilities, we apply CoT fine-tuning and reinforcement learning (RL), inspired by the training strategies of HuatuoGPT-o1 Chen et al. (2024). We leverage a diverse set of high quality financial reasoning datasets—including FinQA Chen et al. (2022), TATQA Zhu et al. (2021), DocMath (Simpshort and Compshort) Zhao et al. (2024b) , DocFinQA Reddy et al. (2024a), and BizBench-QA Krumdick et al. (2024)—covering tasks such as long-context question answering, complex reasoning over financial documents, and table-text integration. We construct high-quality reasoning paths using GPT-4o, combined with a novel data filtering process that selects more challenging samples for training—particularly for the 14B model—ensuring the model learns from difficult, domain-specific reasoning cases rather than generic QA pairs. Experimental results show that this targeted training strategy enables our models to surpass the performance of advanced closed-source models such as GPT-o3-mini and GPT-o1, and approach most powerful LLMs like DeepSeek-V3, DeepSeek-R1, and GPT-4o, demonstrating that domain-specific reasoning-enhanced training with limited, high-quality data can be highly effective.
To the best of our knowledge, this work presents the first comprehensive study of reasoning-enhanced LLMs for financial tasks involving diverse inputs such as financial documents, tables, equations, and structured XBRL texts. First, we conduct a systematic evaluation of 24 general-purpose and reasoning-enhanced LLMs across multiple financial reasoning benchmarks. Second, we introduce Fino1, the first suite of domain-specific, reasoning-enhanced LLMs, trained using GPT-4o-distilled reasoning paths with a novel data curation and filtering strategy. Third, our results show that domain-specific reasoning training with limited high-quality data consistently outperforms general-domain strategies and even surpasses larger closed-source models. We release all models, datasets, and code to support future research in financial AI.
2 Methods
2.1 Datasets
To evaluate the financial reasoning capabilities of reasoning-oriented and general LLMs, we select four diverse benchmark tasks. (1) FINQA Chen et al. (2021) is a large-scale dataset that emphasizes complex numerical reasoning in financial reports, requiring the integration of structured tables and unstructured text, along with a deep understanding of domain-specific terminology. (2) DocMath (Simplong, Complong) Zhao et al. (2024a) evaluates LLMs’ numerical reasoning abilities over long, specialized documents with multi-tiered tables, testing their capacity to extract and reason over extended contexts. The simplong subset contains relatively easier questions, while the complong subset includes more complex questions that require advanced reasoning across tables and context. (3) XBRL-Math Wang et al. (2025) targets reasoning over structured financial data in XBRL filings, which involve standardized taxonomies, hierarchical disclosures, and formulas, challenging models to interpret intricate numerical relationships.

### v3

# Fin-o1: On the Transferability of Reasoning-Enhanced LLMs and Reinforcement Learning to Finance

@@ Abstract
As the fundamental capability behind decision-making in finance, financial reasoning poses distinct challenges for LLMs. Although reinforcement learning (RL) have boosted generic reasoning, the progress in finance is hindered by the absence of empirical study of building effective financial chain-of-thought (CoT) corpus, a systematic comparison of different RL methods, and comprehensive benchmarks. To address these gaps, we introduce FinCoT, the first open high-fidelity CoT corpus for finance, distilled from seven QA datasets by a novel three-stage pipeline that incorporates domain supervision, iterative LLM refinement, and difficulty-aware filtering. Based on FinCoT, we develop Fin-o1, the first open financial reasoning models trained via supervised fine-tuning and GRPO-based RL. Our models outperform existing financial reasoning models and SOTA general models such as GPT-o1, DeepSeek-R1, and GPT-4.5. We also investigate the effectiveness of three different RL methods in improving domain-specific reasoning, offering the first such empirical study. We finally propose FinReason, the first financial reasoning benchmark covering multi-table analysis, long-context reasoning, and equation-based tasks, and evaluate 29 LLMs. Our extensive experiments reveal general reasoning models excel on standard benchmarks yet exhibit obvious performance degradation in financial contexts; even finance-tuned models like Dianjin-R1 and FinR1 degrade on lengthy documents. In contrast, our Fin-o1 models consistently outperform their backbones and larger GPT-o1 and DeepSeek-R1, confirming the effectiveness of our data building and model training strategy. Our study further shows that GRPO yields reliable gains whereas PPO and DPO do not, highlighting the need for targeted data and optimisation rather than scale alone. These findings show that financial reasoning has unique challenges unmet by mere model scaling or generic techniques, highlighting the need for specialized data, models, and optimization. We release all datasets, models, and codes to support future research. 111 Dataset: https://huggingface.co/datasets/TheFinAI/FinCoT
      Fin-o1: https://huggingface.co/TheFinAI/Fin-o1-8B
                https://huggingface.co/TheFinAI/Fin-o1-14B
      Code: https://github.com/The-FinAI/Fino1
@@ 1 Introduction
Financial reasoning is a critical yet underexplored capability for large language models (LLMs), especially as they are increasingly deployed in high-stakes financial applications. Unlike general reasoning tasks, such as mathematical problem solving or code generation, financial reasoning requires models to interpret domain-specific language, analyze structured data like financial tables xie2024open , and extract insights from lengthy, complex documents reddy2024docfinqa . It demands not only numerical computation, but also contextual understanding of financial regulations, economic concepts, and ambiguous textual information xie2023pixiu ; xie2024finben .
Recent advances in reinforcement learning (RL) methods, including proximal policy optimization (PPO) schulman2017proximal , direct preference optimization (DPO) rafailov2023direct , and generalized reinforcement preference optimization (GRPO) tang2024generalized , have substantially enhanced the reasoning capabilities of large language models. These RL-based approaches form the foundation of state-of-the-art (SOTA) reasoning models such as OpenAI’s o1 jaech2024openai and DeepSeek’s R1 guo2025deepseek , achieving notable improvements on mathematical and logical problem-solving benchmarks temsah2024openai ; zhong2024evaluation . By directly optimizing for reasoning performance, RL strategies have become essential for advancing the complex reasoning abilities of LLMs.
However, improving LLMs with RL for financial reasoning is yet hindered by three critical challenges (As shown in Table 1). First, the absence of high-quality, publicly available chain-of-thought (CoT) datasets for financial reasoning limits the ability to model domain-specific reasoning processes. Existing resources are either proprietary jaech2024openai or adapted from general domains guo2025deepseek ; yang2024qwen2 , failing to capture the nuances of financial analysis. No empirical study has asked which source documents or path-generation strategies actually produce effective financial CoT traces. Second, while RL methods have enhanced general reasoning, their effectiveness in financial tasks remains unexplored, with no open-source models demonstrating domain-adapted improvements. The impact of different RL algorithms on financial reasoning remains untested: RL methods like PPO, DPO and GRPO have never been compared side-by-side, leaving practitioners without guidance on the best optimiser for the domain. Third, there is no comprehensive benchmark to systematically evaluate financial reasoning across key challenges such as domain-specific language understanding xie2024open , multi-table reasoning zhao2024docmath , long-context document analysis reddy2024docfinqa , and equation-based reasoning ma2023xbrl . Existing benchmarks like PIXIU xie2023pixiu , FinBen xie2024finben , and BizBench krumdick2024bizbench address only parts of this space. Moreover, current studies only evaluated a few LLM’s financial reasoning ability.
Table 1: Comparison of Existing Reasoning LLMs and Financial Benchmarks. SFT indicates super- vised fine-tuning. RL specifies the reinforcement learning method used. CoT data is categorized into Guide, indicate the presence of gold guidelines to direct reasoning generation, IterR means iteratively verification and refinement, DiffF indicates if involving difficulty-aware filtering to emphasize complex cases. Evaluation indicates whether the benchmarks involve single-table or multi-table reasoning, long-context analysis, and equation-based reasoning. ModN indicates the number of models evaluatd in each work. *Indicates works released after Fin-o1.
Model / Benchmark Domain CoT Data Model Evaluation
Public Guide IterR CompF SFT RL Single Multi Long Context Equations ModN
PIXIU xie2023pixiu Financial $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ 9
FinBen xie2024finben Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ 21
BizBench krumdick2024bizbench Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ 16
DocFinQA reddy2024docfinqa Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ 12
S1-32B muennighoff2025s1 General $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 12
Limo ye2025limo General $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 6
HuatuoGPT-o1 chen2024huatuogpt Other $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ PPO $\times$ $\times$ $\times$ $\times$ 10
DeepSeek-R1 guo2025deepseek General $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ GRPO $\times$ $\times$ $\times$ $\times$ $\times$
OpenAI-o1 jaech2024openai General $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Fin-R1finr1 * Financial $\times$ $\times$ $\times$ $\times$ $\checkmark$ GRPO $\checkmark$ $\times$ $\times$ $\times$ 10
DianJin-R1zhu2025dianjin * Financial $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ GRPO $\checkmark$ $\times$ $\times$ $\times$ 12
Fin-o1 (Ours) Financial $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ GRPO, PPO, DPO $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29
To address these gaps, we present three key contributions. First, we curate FinCoT, the first high-quality, publicly available financial CoT dataset built by a novel three-stage framework that combines domain-aware supervision, iterative LLM refinement, and difficulty-aware filtering. We construct FinCoT from diverse sources including FinQA chen2022finqadatasetnumericalreasoning , ConvFinQA chen2022convfinqa , DocFinQA reddy2024docfinqa , TATQA zhu2021tatqaquestionansweringbenchmark , Econ_Logic quan2024econlogicqa , DocMath-Eval zhao2024docmath , and BizBench-QA krumdick2024bizbench . Our new framework yields high-fidelity financial CoT data that improves 8B and 14B models beyond SOTA reasoning models such as GPT-o1, DeepSeek-R1, and GPT-4.5. Second, we develop Fin-o1 (including Fin-o1-8B and Fin-o1-14) based on FinCoT, the first open-source financial reasoning large language models, trained via a two-phase framework combining supervised fine-tuning and RL-based reasoning enhancement. Fin-o1 consistently outperform existing financial reasoning models and SOTA general-reasoning LLMs such as GPT-o1, DeepSeek-R1, GPT-4.5. We systematically investigate the effectiveness of PPO schulman2017proximal , DPO rafailov2023direct , and GRPO tang2024generalized in improving domain-specific reasoning, offering the first empirical study on RL methods for financial reasoning. Third, we introduce FinReason, the first comprehensive benchmark designed to systematically evaluate the financial reasoning capabilities of LLMs. FinReason covers most LLMs with diverse tasks involving domain-specific language, structured data, and long-context documents, providing a robust framework on datasets such as FinQA chen2021finqa , DM-Simplong zhao2024docmath , XBRL-Math wang2025finnlp , and DM-Complong zhao2024docmath . We conduct a comprehensive evaluation on FinReason, covering 29 large language models from general-purpose, reasoning-enhanced, and financial-specific families, including GPT, LLaMA, DeepSeek, Qwen, and our Fin-o1 models.
Our results reveal that general reasoning models, despite their strong performance on generic benchmarks, struggle with financial reasoning tasks that require precise domain understanding, structured data interpretation, and long-context document reasoning. Frontier financial reasoning LLMs such as Dianjin-R1 and FinR1 exhibit performance degradation compared to their backbones in long-context tasks, highlighting their limited robustness. In contrast, our Fin-o1 models, trained on the FinCoT dataset, consistently outperform their backbone models and surpass larger general-purpose models, demonstrating the importance of domain-specific reasoning supervision. Further analysis shows that reinforcement learning methods have varying effectiveness in this context. While PPO and DPO fail to yield stable improvements, GRPO achieves consistent gains by aligning optimization with task-specific reasoning objectives. These findings illustrate that financial reasoning poses unique challenges that cannot be addressed by scaling model size or applying general reasoning techniques alone, emphasizing the need for specialized data, models, and optimization strategies.
2 FinCoT Reasoning Training Data
In this section, we describe the construction of FinCoT, the first high-fidelity dataset of chain-of-thought reasoning paths designed to enhance and evaluate LLMs’ financial reasoning capabilities, as illustrated in Fig. 1.
2.1 Data Collection
To develop a robust chain-of-thought training dataset tailored for financial reasoning, we systematically curate data from diverse financial data sources. Distinguished from FinR1 finr1 and Dianjin zhu2025dianjin , which only include FinQA for reasoning tasks while the rest of their datasets, i.e, Finance-Instruct flowers2025finance , FinCorpus duxiaoman2023fincorpus , and others that focus primarily on financial knowledge enhancement [^2] Samples of these data can be found in Appendix B.1, our dataset explicitly provides structured chain-of-thought supervision to enable complex financial reasoning, each contributing distinct reasoning skills [^3] See Appendix B for dataset details.:
  • FinQA chen2022finqadatasetnumericalreasoning and ConvFinQA chen2022convfinqa : Focused on numerical reasoning over short financial texts and single well-structured tables, supporting the model in basic arithmetic and aggregation tasks.
  • TATQA zhu2021tatqaquestionansweringbenchmark : Provides semi-structured financial reports with single-table QA tasks, enhancing the model’s ability to reason over less curated, real-world financial tables.
  • DocMath-Eval zhao2024docmath : Adds mathematical reasoning over structured tables, introducing complex quantitative reasoning patterns.
  • Econ-Logic quan2024econlogicqa : Includes financial decision-making and logical inference tasks, requiring models to apply domain-specific logic beyond numerical calculation.
  • BizBench-QA krumdick2024bizbench : Introduces multi-table reasoning tasks, training the model to integrate and synthesize information across multiple data sources.
  • DocFinQA reddy2024docfinqa : Comprises long-form financial documents with QA tasks, enabling the model to handle extended context and multi-step reasoning over lengthy narratives.
Figure 1: Overall framework of development of FinCoT, Fin-o1 and FinReason.
2.2 Iterative Chain-of-Thought Generation and Verification with Domain-Guided Supervision
Building our raw data collection, we further curate the first dataset containing 9,186 chain-of-thought annotated QA pairs, specifically designed to support SFT and RL for financial reasoning models, via a three-stage CoT generation framework by integrating domain-aware supervision, iterative refinement, and difficulty-aware filtering. This framework addresses the limitations of existing chain-of-thought methods, which prompt LLMs such as DeepSeek-R1 or Gemini team2024gemini ; ye2025limo ; muennighoff2025s1 to generate reasoning paths in a single round finr1 ; zhu2025dianjin , sufficient for general tasks but inadequate for financial reasoning which requires precise, multi-step, domain-informed calculations. For example, financial tasks like calculating profits require carefully following each calculation step with the right data and rules, where even small mistakes can completely invalidate the final answer.
In the first stage, we tackle the lack of structured guidance in existing methods by incorporating programmatic steps and expert-curated sub-questions from datasets such as BizBench and ConvFinQA, collected during our data preparation phase. While ConvFinQA offers high-quality sub-questions, it does not provide complete question-answer pairs that capture the full financial decision-making process. To close this gap, we use GPT-4o to synthesize these sub-questions into combined, holistic questions (see Fig. 2. This allows the model to see both the detailed steps and how they fit into a broader financial reasoning workflow, helping it generate more coherent and logically consistent reasoning paths that mirror domain expert expectations).
Figure 2: Workflow of curating the combined question and examples of program progress.
In the second stage, we address the brittleness of single-pass reasoning by introducing an iterative verification and refinement process, enabling the model to correct its reasoning progressively. Inspired by HuatuoGPT-o1 chen2024huatuogpt but adapted to the stricter demands of financial reasoning, we prompt GPT-4o to generate step-by-step reasoning paths, which are independently evaluated by a verifier (also GPT-4o). Instead of discarding incorrect outputs entirely, our framework forces the model to revisit earlier steps, refine its assumptions, and adjust its reasoning flow based on verifier feedback (see Fig. 3 in appendix C.). This allows the model to detect and fix errors that might otherwise propagate unchecked, ensuring greater accuracy and robustness, especially in complex financial tasks where early missteps can have outsized impacts.
In the third stage, we introduce a difficulty-aware filtering step to ensure that the final dataset targets genuinely challenging reasoning scenarios (as shown in Fig. 1). Unlike prior methods chen2024huatuogpt ; finr1 that focus only on verifying correctness, we also consider question difficulty, motivated by recent studies showing that easy QA pairs do little to advance reasoning ability ye2025limo ; muennighoff2025s1 . We use Llama3.1-8B-Instruct to attempt each question, then apply a verifier (see Section 4.3) to check the answers. QA pairs that are correctly solved by the filtering LLM are excluded, which is deemed less challenging.
3 Fin-o1
In this section, we propose Fin-o1 (including Fin-o1-8B and Fin-o1-14B) based on FinCoT, the first open source LLMs targeted on financial reasoning tasks. Qwen3 are selected as backbone models, given their strong performance among small and mid-sized LLMs (shown in Table 3). Following previous studies chen2024huatuogpt ; dubey2024llama , a two-stage training approach, supervised fine-tuning and RL, is adopted to develop the model. FinCoT is divided into two parts FinCoT_SFT (7686) and FinCoT_RL (1,500) to support the training. Based on FinCoT, we further conduct the first evaluation of different RL methods to assess their impact on financial reasoning tasks.
3.1 Supervised fine-tuning
In the first stage, we perform supervised fine-tuning, with the FinCoT_SFT, aiming to guide the model to systematically structure its reasoning process before providing the final answer. As shown in the right part of Fig. 1, each training instance is composed of a prompt, contextual information extracted from textual and tabular data, and a question that requires multi-step reasoning. The output target consists of a detailed reasoning path, followed by a summarized answer, emphasizing the importance of step-by-step logical analysis before deriving the final response. By structuring the output in this manner, the objective of SFT is to cultivate a disciplined reasoning framework, encouraging the model to first articulate a coherent analytical process, thereby enhancing interpretability and accuracy in complex financial reasoning scenarios. After the supervised fine-tuning, we get Fin-o1-SFT.
3.2 Reinforcement learning
In the second stage, we conduct a comprehensive evaluation of various RL methods to assess their impact on financial reasoning tasks. To effectively evaluate these methods, we specially design targeted reward functions and curate high-quality positive and negative samples tailored to each method. This comparative analysis aims to provide insights into the strengths and limitations of each RL method in enhancing reasoning capabilities for complex financial tasks.
Proximal policy optimization.
We start from PPO schulman2017proximal . It optimizes the policy by maximizing the expected reward while constraining the extent of policy updates, thus preventing excessive divergence from the reference policy. Each input $x$ is processed by the Fin-o1-SFT to generate a response with a reasoning path $y^{*}$. The reward $r$ is defined based on the verifier feedback when comparing if the $y^{*}$ is consistent with the gold answer $y$. To effectively evaluate the generated answer with the gold, GPT-4o is used to verify the correctness of the results.
$r(x,y,y^{*})=\begin{cases}1&\text{if }\text{verifier}(y,y^{*})=\text{True}\\ 0&\text{otherwise}\end{cases}$ (1)
Additionally, the reward is adjusted by incorporating the Kullback-Leibler (KL) divergence penalty between the updated policy $\pi_{\theta}$ and the reference policy $\pi_{ref}$ luong2024reft , aiming to prevent excessive deviation from the reference policy while ensuring stable learning.:
$r^{\prime}(x,y,y^{*})=r(x,y,y^{*})-\beta\cdot KL(\pi_{\theta}\parallel\pi_{ref})$ (2)
Direct preference optimization.
DPO rafailov2023direct refines the model’s decision-making by systematically contrasting positive and negative samples through pairwise comparisons. In our framework, we extend the conventional DPO setup by implementing a more strategically designed sampling mechanism tailored to the nuances of complex financial reasoning paths.
The iterative CoT reasoning path generation process is leveraged in this step to effectively curate high-quality positive-negative samples. As detailed in 2.2, this process involves employing GPT-4o to generate reasoning paths that undergo iterative verification and correction. Rather than employing low-quality negative samples that diverge significantly from the positive samples, we implement a targeted selection strategy that designates the last failed reasoning path $R_{n-1}$ as the negative sample. This path is specifically chosen as it closely mirrors the input query in logical structure while containing identifiable reasoning flaws, making it a more informative negative sample. By focusing on plausibility and logical coherence, our sampling strategy mitigates the risk of selecting trivial or overly simplistic negative samples, thereby reinforcing the model’s capacity to differentiate between nuanced reasoning paths and effectively learn from more challenging and contextually relevant negative samples.
Generalized proximal policy optimization.
Compared with PPO that optimizes the model by maximizing expected rewards through constrained policy updates, GRPO tang2024generalized extends PPO by incorporating additional objectives or multi-task learning capabilities, making it suited for financial reasoning tasks where accuracy and reasoning completeness, long-context reasoning abilities are crucial for effectively handling complex financial scenarios.
Therefore, to comprehensively improve LLMs’ reasoning abilities in financial reasoning tasks, we propose a novel multi-faceted reward structure in GRPO that systematically evaluates four critical dimensions. Compared with prior work that predominantly targets accuracy and format chen2024huatuogpt ; finr1 ; zhu2025dianjin , our framework introduces two additional components: reasoning logic and length rewards. By integrating these dimensions, we aim to mitigate shortcut learning, reinforce logical consistency, and incentivize effective reasoning over long contexts—a crucial capability given the inherently lengthy and complex nature of financial documents. Formally, the reward function for GRPO is designed as follows:
$r(x,y,y^{*})=\alpha_{1}\cdot r_{acc}+\alpha_{2}\cdot r_{logic}+\alpha_{3}\cdot r_{format}+\alpha_{4}\cdot r_{length}\cdot r_{acc}$ (3)
where:
  • $r_{acc}$: Accuracy reward, based on the verifier’s assessment of the final response correctness.
  • $r_{logic}$: Logical consistency reward, evaluating the correctness of the reasoning path.
  • $r_{format}$: Assessing if the generated response follows the designed format.
  • $r_{length}$: To encourage the model’s ability to reason over lengthy contexts—a critical requirement in financial analysis given the extensive nature of many financial documents—we provide an additional reward when the answer is correct and the context length exceeds 8,192 tokens. If the accuracy reward is zero, the length reward is also set to zero.
To effectively evaluate the generated response, we also leverage LLMs (GPT-4o) as judge. Following the reward score design in finr1 ; chen2024huatuogpt , we first assesses the accuracy of the generated response, signing a reward of 1 if the answer is deemed correct. Similarly, if the logical reasoning is validated as correct, a reward of 1 is also given. For format and length evaluation, we employ rule-based methods to assign corresponding rewards, the format reward is set as 0.1 and the length reward is set as 1.
4 FinReason Evaluation Benchmark
In this section, we present FinReason, the first comprehensive benchmark listed in Table 2 designed to systematically evaluate the financial reasoning capabilities of LLMs.
4.1 Tasks
Financial reasoning tasks present distinct challenges beyond general reasoning, as they require not only understanding the structure and semantic content of financial documents and tables but also accurately interpreting and executing complex equations. To this end, we introduce a unified benchmark that systematically integrates reasoning over (1) financial texts with tables, (2) long-context reasoning, (3) multi-table analysis, and (4) interpretation and execution of financial equations, establishing a comprehensive evaluation framework to assess and elevate LLM capabilities in financial reasoning tasks. Four datasets are selected to evaluate LLMs across these key aspects of financial reasoning:
(1) FinQA chen2021finqa serves as the foundational benchmark for quantitative reasoning over structured tables and basic financial texts. It assesses the model’s ability to extract and synthesize financial information related to core financial concepts such as revenue, profit margins, and financial ratios.
(2) Simplong zhao2024docmath targets long-context reasoning within single-table scenarios. It assesses the model’s capacity to perform numerical reasoning based on structured data while maintaining context throughout extended financial texts.
(3) Complong zhao2024docmath extends Simplong by introducing multi-table scenarios, challenging the model to integrate information across multiple tables. This dataset tests the model’s ability to handle more complex financial reporting contexts, where financial data is fragmented and requires cross-table reasoning to draw accurate conclusions.
(4) XBRL-Math wang2025finnlp emphasizes equation-based reasoning, requiring the model to accurately interpret and execute financial formulas derived from structured data in XBRL filings. This dataset assesses the model’s ability to associate financial terms with corresponding numerical data and perform structured calculations.
Table 2: Overview of the datasets evaluated in the study.